UT Dec 11 2022 - Split Night PETS/DEC



Night Info

  • Sunset 17:19
  • Nautical Twilight Ends 18:12
  • Astro Twilight Ends 18:42
  • Moonrise 19:15
  • Astro Twilight Starts 05:44
  • Nautical Twilight Starts 06:14
  • Sunrise 07:12
  • Moon 94% illumination

Staff

  • OSA: S.Allanson
  • ISp: Mtn
  • Mtn Manager: P.Hartley
  • ISA/SO: O.Kuhn

Night Plan

  • Staring Config PEPSI-Bino
  • PEPSI-PETS Twilight-03:37
  • Reconfigure PEPSI-PFU to MODS 03:38-03:50

MODS Checkout

  • Highest priority (figuring a total of ~1.75-2 hrs to go through high+med priority tasks)
  • /home/lbto/Service_Observing/2022B_D_E/MODS/11Dec22
  • AGw checkout (high pri; ~30 min to 1 hr)
    • Check WFS hotspot
    • Check alignment of rotator center on guider to WFS hotspot
  • RotCen: Rotator Center on MODS (med-high; 10 min): M52, Hyades, Coma Star Cluster (use pointing mask)
  • RZP: Rotator Zero Point (high; 15-20 min; tonight just one mask field)
    • masks for 00h25d, 04h25d,*06h80d* and 08h25d to cover night
    • g06p80 and g08p25 avail from handover (10:38 UT) until twilight, g06p80 is 56 deg from moon but g08p25 just 1.5 deg.
    • From 20220925 wiki: Rotational offsets needed to align MODS MOS masks (2021B to 2022A). There were 6 more acquisitions of 2 masks (UDS1 and UDS2) done during partner science 10-17-10-21 UT and all of these require a -0.15 to -0.17 deg rotational offset, both for MODS1 and MODS2 (with the exception of one MODS1 acquisition which required only -0.02 deg). These are UDS1 ID503205 & UDS2 ID593032. Because it is both MODS1 & MODS2...curious about the astrometry on the image(s) used to make the mask.

instr mean median sig_n-1 n
MODS1 -0.03 -0.03 0.05 33
MODS2 -0.05 -0.04 0.05 55
  • Spectrophotometric Standard (high; 35-min, 15(acq)+20min(obs 1x per mode)): 1spec/mode mode if non-phot, 3 spec/mode if phot
    • Feige34 AM > 1.09 from 10:50 to twilight, 38 deg from moon. 50-min for 3x all modes, but 20 min for 1x all modes.
    • G191-B2B AM 1.43-1.74 bt 11-12 UT, 41 deg from moon
  • Photometric imaging field (if phot, high; 24-min for dual+dir modes: 3x 8-10 min/mode)
    • 11h60d sdss field: 50 deg from moon and with ~half-dozen stars with ugriz photometry from DR16. 60-sec per image, 2x @ u.
    • 10.5h00d sdss field: 46 deg from moon, in case we cannot observe to NE
  • BlindOffset: long-slit alignment and blind offset check on mask field with 2.4" slit (low; 20-min)
    • r06d80_a2b.txt and r06d80_c2d.txt
  • Dithering Imaging Scripts (low; 10-min): on mask field with radec and detxy offsets (this is to see if we run into any offset errors)
    • 11h60d sdss field at g and r with +/-35" dithers.

LUCI-AO/ESM/SuperESM testing

  • If MODS tasks are completed - then go for this task until twilight - won't be able to do all or most targets - order of priority (or depending on winds)
  • N375 w/Full AO -2.5hr
    • Conditions: seeing 1.2" or better, thin clouds OK
    • OT: /home/lbto/Service_Observing/2022B_D_E/Rothberg/DwarfMerger/Dwarf_Mergers_V3.xml
      • SDSSJ1148+2149 - Bin4 Full AO, SL in parallel - visible 02:30-twilight
        • Script (pre-generated) in SuperESM folder: SDSSJ1148+2149_AO_SL.xml
      • SDSSJ1307+5425 - Bin3 Full AO, SL in parallel - visible 02:30-twilight
        • Script (pre-generated) in SuperESM folder: SDSSJ1307+5425_AO_SL.xml
    • OT: /home/lbto/Service_Observing/2022B_D_E/LUCI/AO/AO_Testing/AO_Testing_V111.xml
      • FF1106+32-1 - Bin 4 Full AO, SL in parallel - visible 01:15-twilight - already observed with ESM
        • Script (pre-generated) in FullAO _N375: FF1106+3201_AO_SL.xml
    • If time:
      • SDSSJ0829+1427 - Bin2 Full AO, SL in parallel - visible 23:30-twilight
        • Script (pre-generated) in SuperESM folder: SDSSJ0829+1427_AO_SL.xml
      • SDSSJ1025+4848 - Bin3 Full AO, SL in parallel - visible 00:30-twilight
        • Script (pre-generated) in SuperESM folder: SDSSJ1025+4848_AO_SL.xml
      • the folder also has _L1.xml versions if LUCI-2 is not available
      • flats (L1 and Bino) in SuperESM folder


Poor Weather Backup

  • Local Mergers - bino LUCI
    • OT File: /home/lbto/Service_Observing/2022B_D_E/Rothberg/LocalMergers/COBandhead_Mergers_V57.xml
    • Finder Charts: /home/lbto/Service_Observing/2022B_D_E/Rothberg/LocalMergers/finder_charts
    • Any conditions - go longer if really bad seeing and/or thick clouds
    • Any conditions - go longer if really bad seeing and/or thick clouds
    • Mono version available if a LUCI goes down
    • Run these BEFORE moving to the ICRF3 backup program
    • Targets
      • NGC 2744 - RA 09hr Dec +18 , guidestar R=12.6 - visible 00:00-twilight
  • ICRF3 - bino LUCI
    • OT File: /home/lbto/Service_Observing/2022B_D_E/Rothberg/ICRF3/ICRF3_LUCI_V31_LBTO_AZ.xml
    • Finder Charts: /home/lbto/Service_Observing/2022B_D_E/Rothberg/ICRF3/finder_charts
    • Any conditions
    • Should be sufficient targets to cover entire part of the night
    • LUCI-1 zJspec G200 1" slit, LUCI-2 HKspec G200 1"slit (all but one)
    • if only one LUCI available scripts is zJspec
    • Pick Targets in OT based on visibility (anything RA>04hr) and best to avoid clouds, winds, etc. Target name has RA and Dec in it (e.g. IERS 0509+293 is RA 05hr Dec +29)

Night Log

00:12 Steve opened the enclosure: Starting with bino PEPSI-PFU. Clear skies. Temp 1C, RH 20%, winds gusting to 6m/s from the SW

00:59 Woke up MODS and ran the simSnap script (while Steve was collimating near first PEPSI target).

01:08 Launched luci_obs_test (test version that Glenn described in his email from 6Dec22). Initialized both LUCIs and switched to the N375 camera and no mask in both. Checked alignment with N375/nomask. OK on L1 and quite good on L2 - minor adjustment of +5,-22 pix on L2.

01:32 At 28 deg elev, T=0.5 deg C, there are problems with mirror limits (some question about whether latest change to thermal coeffs in lookup tables are ok). Steve called J Hill, but got the telescope collimated so observations are proceeding. (Object didn't rise above 26 deg until around now, before that observed Io).

10:19 So far, so good with the PETS transit. Seeing has been superb, <0.5", for most of it, although now it is climbing to ~1".

10:38 Done with the transit. Reconfiguring to MODS. We noticed an oddity with the LBTplot/mag and PEPSI/flux plots around 10:10 UT - the SX/DX flux/mag signals dropped although there was not a new preset, we were tracking the same star for hours. When Steve changed the guide camera exptime, the flux/mag dropped although these should be normalized to 1-sec.

AGw checkout

10:50 Guider hotspot. I relaxed (+2) the values for maxOffsetToHotspot and maxHotspotFromProbe in MODS_(L,R).cfg and Steve bounced GCS.

11:00 Steve is pointing and collimating. Colimating. The blue circle is huge. The guider hotspot looks fine.

(Note added later: Before the next task (align RC to hotspot) I restored maxOffsetToHotspot and maxHotspotFromProbe to their nominal values in MODS_(L,R).cfg and Steve bounced GCS. The larger than normal values explain why the blue circle on GCS was huge.)

11:14 Align rotator center to hotspot. On this preset, the blue circle shrunk to correct size.
  • For DX center of rotation is 187.41, 161.45 (/lbt/data/share/tcs/lbto.20221211/right_guider000001.fits). Add to that 84, 50 (x,y crop) to get 271.41, 211.45. GCS hotspot is currently: 267, 240. Y is off - known.
  • For SX center of rotation is 177.33, 173.31 (/lbt/data/share/tcs/lbto.20221211/left_guider000004.fits). Add +75, +21 gives 252.33, 194.31. GCS hotspot is currently 254.8, 201.0
  • Steve confirmed that the ptincrements I was giving did not make to the mirrors. We may want to consider the order of operations for this procedure (does the ptincrement have to be to both sides simultaneously? Does it have to be done when the preset is functional, i.e. before moving the rotator?)

Rotator center on detector

11:45 Rotator center on detector
  • Active preset to Coma Star Cluster field
PA=0 active 3 3 3 3
PA=180 track 4 4 4 4
11:55 Steve is correcting the pointing. There was a large shift between the active and track presets, on both MODS (Note: this was not a shift, but I jumped ahead to run the PA180 track preset, so it was just that the field had rotated). Also, I'm noting that the MODS2 images are not well collimated. (Note: The sieve spots are not perfectly round and they are a bit larger on MODS2 B & R than on MODS1, but I don't think the defocus looks bad enough to cause the elongated sky images that we are seeing).

12:01 Active PA=0 to collimate (don't take images).
PA=0 active - - - -
PA=0 track 5 5 5 5
PA=180 track 6 6 6 6

12:03 Doug starting W unit on LUCI1.

Rotator zeropoint check

12:12 Rotator zeropoint check: acqBinoMODS g06p80ep23acq.txt -- guide star was a close double, both components were in the lenslet. Chose a new guide star and resent the script.

m1r: 8 & 9 --> offsetpointing 0.0086 -0.128 0.743 detxy rel

m2r: 8 & 9 --> offsetpointing -0.0176 4.196 -1.355 detxy rel

Spectrophotometric Standard

12:33 Standard Star: acqBinoMODS feige34.acq (the star came up well-centered in the slit in MODS1, and modsAlign was only making it go back and forth).

m1r: 11 --> offsetxy -0.467 10.046 rel --> 12 --> offsetxy 0.455 0.247 rel --> 13 --> offsetxy -0.690 -0.014 rel --> 14, 15 --> offsetxy 0.700 -0.205 rel --> 16, ok

m2r: 11 --> offsetxy 4.471 7.016 rel --> 12 --> offsetxy -0.471 -0.261 rel --> 13 --> offsetxy -0.314 -0.163 rel --> 14, 15 --> offsetxy 0.565 -0.357 rel --> 16, ok

Seeing 0.9, 1.1 on SX/DX but changing

12:50 execBinoMODS feige34_1x_all.obs (probably will stop it before we get to the prism).
  m1b m1r m2b m2r
dual 7 17 7 17
red   18   18
blue 8   8  
Seeing 0.7/0.8" and airmass 1.02

Doug is shutting down the OCAM & signing off... unfortunately we have not had any time for the superESM tests.

Photometric imaging field

13:06 Photometric Imaging standards: 10h30m +00d field
  • a few timeouts on the mods1 instconfig
  • MODS2 red image is very elongated (19) MODS2 blue is also elongated, though not exactly the same elongation & PA.
  • m2b: 10 & 11, m2r: 19, 20, 21
  • m1b: 10 & 11, m1r: 19, 20, 21

13:21 Direct Mode blue then red launched on MODS2
  • m2b: 12 & 13, m2r: 22,23,24

13:25 Direct Mode launched on MODS1

13:28 It's getting very light...

13:37 MODS1 i-band image is saturated, z-band stopped early.

Twilight sky flats

13:39 Slewing to Blank 09-07 for skyflats. dual g & z, dual u & i and 2-3 blue-only u band

14:01 Steve is closing up.

14:11 Taking the remaining calibrations (Blue-Only Pixel flats)
  m1b m2b
blue pixflats 24- 23-

Summary

The highest priority MODS checkout tasks were completed and the instrument seems ready, but with a caveat that the image quality on MODS2 doesn't look good. I obtained images in dual, red and blue only modes so I will check whether this is seen in all modes or just some. I checked the sieveMask images and, while not perfectly round and the FWHM of the spots near the center is a bit larger on MODS2 (0.28-0.29") than on MODS1 (0.25"), I don't think this can account for the elongation seen in images such as mods2r.20221211.0019.fits, mods2b.20221211.0010.fits or mods2*.2022121.0003.fits. The AGw checkout gave results similar to before: the offset between the rotator center on the guide camera & the hotspot is < 8 pix except for the Y coordinate for MODS2, and that is a known issue. The rotator zeropoints are consistent with the current transforms. I obtained spectra of the spectrophotometric standard Feige 34 in all grating modes, and images of a field with photometry from SDSS DR16 in dual mode; I tried direct modes but it was already getting very light by that time.

-- BarryRothberg - 08 Dec 2022
Topic revision: r11 - 11 Dec 2022, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback